A cloud has a potential difference relative to a tree of 6.8 x 102 MV. During a lightning storm, a charge of 1.60 x 102 C travels through this potential difference. How much work is done on this charge?
1
Expert's answer
2020-02-27T10:04:45-0500

The work done by an electric field


"W=qV""=1.60\times 10^2\:\rm C\times 6.8\times 10^8\:\rm V=10.88\times 10^{10}\:\rm J."